OpenStudy (fanduekisses):

How do you integrate x^(7/2) ln x dx ?

OpenStudy (fanduekisses):

\[\int\limits_{}^{}x^{7/2}lnx dx\]

OpenStudy (fanduekisses):

u= x^7/2 dv=lnx v=? idk lol i'm stuck du= 7/2x^5/2

OpenStudy (sshayer):

\[\int\limits x ^{\frac{ 7 }{ 2 }}\ln x~dx=\ln x \frac{ x ^{\frac{ 9 }{ 2 }} }{ \frac{ 9 }{ 2 } }-\int\limits \frac{ 1 }{ x }* \frac{ x ^{\frac{ 9 }{ 2 }} }{ {\frac{ 9 }{ 2 }} }dx\] =?

OpenStudy (sshayer):

integration by parts

OpenStudy (sshayer):

ln x is first function

zepdrix (zepdrix):

ya you've got your parts backwards :) remember that when you differentiate natural log, it turns into some of that x business. So you want your ln x = u

OpenStudy (sshayer):

\[\int\limits uv dx=u \int\limits vdx-u' \int\limits vdx\]

OpenStudy (agent0smith):

"u= x^7/2 dv=lnx v=? idk lol i'm stuck" ^Right there you should have known that you chose u and dv incorrectly.